OpenStudy (anonymous):

Nick had 67.50 in pennies and dimes. If there were 150 more pennies than dimes, how many dimes dos he have?

OpenStudy (anonymous):

let n for nick=$.05, p for a penny=.01 .01p+.05n=$67.50 p=150+n Subtitution: p=150+n .0.1(150+n)+.05n=$67.50

OpenStudy (anonymous):

can you do from now?

OpenStudy (anonymous):

very easy

OpenStudy (anonymous):

yes thank u

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

thank you for medal

OpenStudy (anonymous):

sorry I look wrong! penny and dime not penny nick let d for dime=$0.10d p for a penny=.01p .01p+.10d=$67.50 p=150+d Subtitution:p=150+n .01(150+d)+.10d=$67.50 $1.5+.01d + .10d=$67.50 0.11 d =$67.50 -$1.5 d=$66/0.11 d=600cents Subtitution: d=600 .01p+.10(600)=$67.50 .01p+.10(600)=$67.50

OpenStudy (anonymous):

.01p+.10(600)=$67.50 .01p +60 =$67.50 .01p =$67.50-60 p= 7.5/0.1 p= 750 cents check: .01(750)+.10(600)=$67.50 $7.5 + $60.00 = $67.50 Nick have : $7.5 or 750 penny 600 cents or $60.00 Dime